how much must a man 35 years of age deposit in a savings account today in a bank paying 4 1/2 % compounded quarterly in order to have $25000 in the account 25 years from now?
Answer by stanbon(75887)

how much must a man 35 years of age deposit in a savings account today in a bank paying 4 1/2 % compounded quarterly in order to have $25000 in the account 25 years from now?
----
A(t) = P(1+(r/n))^(nt)
----
25000 = P(1+(0.045/4))^(4*25)
---
250000 = P(3.0609)
P = $81674.51
==================
Cheers,
Stan H.
====================